OpenStudy (anonymous):

A ball is dropped from a height of 3 ft. The elasticity of the ball is such that it always bounces up one-third the distance it has fallen. (a) Find the total distance the ball has traveled at the instant it hits the ground the fourth time. (Enter your answer as an improper fraction.)

OpenStudy (anonymous):

is the answer 40/9?

OpenStudy (anonymous):

no that doesnt work

OpenStudy (anonymous):

This how I solved it: Suppose K=1/3 first time d1=3 second time d2=3*K=1 Third time d3=1*k=1/3 Forth time d4=K(1/3)=1/9 Now adding up d1+d2+d3+d4, we get the total distance d=40/9

OpenStudy (anonymous):

Sorry I made a huge mistake. Im gonna correct it

OpenStudy (anonymous):

That looks right to me but my online homework says its wrong...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Suppose K=1/3 first time d1=3 second time d2=2*(3*K)=2 Third time d3=2*(2*k)=4/3 Forth time d4=2*K(4/3)=8/9 Now adding up d1+d2+d3+d4, we get the total distance d=65/9. Is that correct. Hopefully yes

OpenStudy (anonymous):

it still says its wrong...?

OpenStudy (anonymous):

What is the correct answer?

OpenStudy (anonymous):

IDK. in my book it gives a similar problem who's answer is 17 8/9

OpenStudy (anonymous):

its the same problem only the ball is dropped from 9 feet instead of 3 ft

OpenStudy (anonymous):

What you mean by 17 8/9?

OpenStudy (anonymous):

the answer is a mixed number

OpenStudy (anonymous):

Convert this number in a decimal number. Cannot remember how to do that!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is 17(8/9) the answer for the problem you have to solve? It doesnt make any sense!

OpenStudy (anonymous):

no that was the answer to a similar problem. I thought if i gave you that answer then it might help you figure out the the problem

OpenStudy (anonymous):

Yeah coz im trying to help you. Or are you challenging me?

OpenStudy (anonymous):

A ball is dropped from a height of 9 ft. The elasticity of the ball is such that it always bounces up one-third the distance it has fallen. (a) Find the total distance the ball has traveled at the instant it hits the ground the fourth time. the answer to this is 17 (8/9)

OpenStudy (anonymous):

But I need the answer for the problem your are solving now dude.

OpenStudy (anonymous):

I dont know the answer to that one thats why im on here!

OpenStudy (anonymous):

I got 53/9

OpenStudy (anonymous):

Convert 65/9 in a mixed number

OpenStudy (anonymous):

53/9 works!

OpenStudy (anonymous):

Convert 211/27 ins a mixed number too

OpenStudy (anonymous):

What formula did you use?

OpenStudy (anonymous):

Can you write down step by step, how you got 53/9?

OpenStudy (anonymous):

You start at 3 fall 3 rise (3 * 1.3) =1 fall 1 rise ( 1 * 1/3) = 1/3 fall 1/3 rise (1/3 * 1/3) = 1/9 fall 1/9 I stop at 1/9 because at fall 1/9 is the 4ht instant I hit the ground. add 3 + 1 + 1 + 1/3 + 1/3 + 1/9 + 1/9 5+ 2/3 + 2/9 5 + 6/9 + 2/9 5 + 8/9 45/9 + 8/9 53/9

OpenStudy (anonymous):

The formula of the the that is droped from 9 ft is 18-(1/3)^n-3. so i assume that the ball droped from 3ft had a similar formula...

OpenStudy (anonymous):

the= ball

OpenStudy (anonymous):

I can derive a formula but it will be hard since you start from the top and you want to know the distance it travels on the fourth hit to the floor.

OpenStudy (anonymous):

ok thats cool. you got the answer right. Thanks!

OpenStudy (anonymous):

btw the formula you use gives you 53/3 not 53/9 for the 4th hit.

OpenStudy (anonymous):

Are you talking about the problem were the ball is droped from 9ft?

OpenStudy (anonymous):

THe way I visualized it was right. However I made a mistake in fraction reduction. Sorry to dissapoint

OpenStudy (anonymous):

Hey its cool man. I appreciate the help!

OpenStudy (anonymous):

OK thanks. I like to help